# Minimisation of Laplacian eigenvalue with indefinite weight under inhomogeneous Robin boundary condition

### v1

@@ Abstract
This paper explores a certain Laplacian eigenvalue optimisation problem with indefinite weight under inhomogeneous Robin boundary condition. The minimum principal eigenvalue is fully determined in one dimension by formulating the problem as a shape optimisation one. The result is verified numerically using a shooting method.
pacs:
MSC Classification]35P15, 35J25
@@ 1 Introduction
In this paper, we consider the following eigenvalue problem
$\left\{\begin{aligned} \Delta\phi+\lambda m\phi&=0&\ \text{in}&\ \Omega\\ \partial_{\mathbf{n}}\phi+\beta\phi&=0&\ \text{on}&\ \partial\Omega\end{aligned}\right$. (1)
where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$ with Lipschitz boundary, $\mathbf{n}$ is the unit outer normal on $\partial\Omega$, and the weight function $m$ is a bounded measurable function that changes sign in $\Omega$, i.e. $\Omega_{m}^{+}:=\{x\in\Omega;\ m(x)>0\}$ satisfies $|\Omega_{m}^{+}|\in(0,|\Omega|)$, and that
$1\leq m(x)\leq\kappa\ \text{a.e.}\ x\in\Omega$
for a constant $\kappa>0$. The inhomogeneous Robin boundary condition applies, with $\beta:\partial\Omega\to\mathbb{R}$ a piecewise continuous function.
According to [1], [2] and [3], $\lambda$ is called a principal eigenvalue of (1), if the associated eigenfunction $\phi$ is positive, where $\phi\in H^{1}(\Omega)$.
The investigation of Laplacian eigenvalues with an indefinite weight under Robin boundary conditions originates from studies of diffusive logistics models in a bounded domain $\Omega$ with suitable boundary. In [4], it is shown that the density $u(x,t)$ of the species at location $x$ and time $t$ satisfies the following equation:
@@
$\begin{cases}u_{t}=\Delta u+\omega u[m(x)-u]\ \text{in}\ \Omega\times\mathbb{R}^{+},\\ \partial_{\mathbf{n}}u+\beta_{0}u=0\ \text{on}\ \partial\Omega\times\mathbb{R}^{+},\\ u(x,0)\geq 0,\ u(x,0)\not\equiv 0\ \text{in}\ \bar{\Omega},\end{cases}$ (2)
where $\omega>0$ is a certain parameter. The Robin boundary parameter $\beta_{0}$ represents the inhospitableness of $\partial\Omega$. If $\beta_{0}=0$, the boundary is a barrier; if $\beta_{0}=+\infty$, the boundary is hostile. $m$ represents the intrinsic growth rate of the species, changing sign in $\Omega$, so that $\int_{\Omega}m$ represents the total resources in the spatially heterogeneous environment. In [5], the relation between the minimal positive principal eigenvalue $\lambda=\lambda(m)$ of (1) and the parameter $\omega$ in (2) is discovered: if $\omega\leq\lambda(m)$, then $u(x,t)\to 0$ uniformly in $\bar{\Omega}$ as $t\to\infty$ for proper initial data, i.e. the species extinguish; if $\omega>\lambda(m)$, then $u(x,t)\to u^{*}(x)$ uniformly in $\bar{\Omega}$ as $t\to\infty$, where $u^{*}$ is the unique positive steady solution of (2) in $W^{2,q}(\Omega)$ for every $q>1$, i.e. the species survives. Consequently, it is natural to consider the problem of minimising such an eigenvalue.
In this paper we consider, more generally, the inhomogeneous Robin boundary condition, and we focus on the case $\beta\geq 0$ on $\partial\Omega$.
@@
We first prove the existence and uniqueness of a positive principal eigenvalue following the method by [6].
Lemma 1 ([7], Theorem 11.4 and 11.10).
@@
the map $\mu:\lambda\mapsto\mu(\lambda)$ is also concave as the infimum of a collection of concave functions. In other words, $\mu^{\prime\prime}(\lambda)\leq 0$. Meanwhile, it is obvious that there exist $\phi_{1},\phi_{2}\in H^{1}(\Omega)$ satisfying $\int_{\Omega}m\phi_{1}^{2}>0$ and $\int_{\Omega}m\phi_{2}^{2}<0$, hence $\lim_{\lambda\to+\infty}\mu(\lambda)=\lim_{\lambda\to-\infty}\mu(\lambda)=$-$\infty$. Thus, $\mu=\mu(\lambda)$ increases until it attains its maximum, and decreases thereafter.
@@
and thus $\mu(0)=0$. Since the associated eigenfunction is clearly a constant, by Lemma 2, we get:
@@
  • if $\int_{\Omega}m>0$, then $\mu^{\prime}(0)=0$, and $\mu$ has no other zeros, hence (1) has a unique principal eigenvalue 0.
@@ 2 The 1-dimensional case
In this section, we derive the complete solution of the problem (1) in one dimension.
@@
Now let the optimal set be the interval $E=(a,b)$ where $0<a<b<1$ and c=bac=b-a. We intend to find the optimal value of $a$ following the method by [2]. Let $\beta_{0}=\beta(0)\geq 0$ and $\beta_{1}=\beta(1)>0$. Rewrite the problem using $m=\kappa\chi_{E}-\chi_{\Omega\setminus E}$, and we have
@@
where $C_{i}=C_{i}(E)$ are constants for $1\leq i\leq 6$. Now we simplify the boundary and transmission conditions in (6). By organising the coefficients of $C_{i}$ ($1\leq i\leq 6$), we get
@@
First we order the term and get
@@
Let us first verify (11) in two extreme cases. When $\beta_{0}=\beta_{1}=0$, the problem is reduced to Neumann boundary condition, and we have obtained the characteristic equation [^2] We remark that the characteristic equations (18) and (20) obtained in [2] based on homogeneous Robin boundary condition admit the same typo on the left-hand side.
@@
On the other hand, when $\beta_{0},\beta_{1}\to+\infty$, the problem is reduced to Dirichlet boundary condition, we similarly get the characteristic equation
@@
Now we use the product-to-sum formulae for hyperbolic functions to simplify it. Using $a\sqrt{\lambda}+\sqrt{\lambda}(1-b)=\sqrt{\lambda}(1-c)$ we get
@@
It is shown in Theorem 4 that $\partial_{\beta_{0}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$ and $\partial_{\beta_{1}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$. Now we have the implications:
@@
We prove in Lemma 7 that $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)$ cannot vanish for any $a\in[0,1-c]$. Therefore, for each $a$, it has a unique zero
@@
Denote by $h=h(c,\kappa,\lambda)$ the expression on the right-hand side of (39). We show in Lemma 8 that $h<1$ for any $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$ under the assumed constraints on $c$, so $A(a)<0$ for any $a\in[0,1-c]$. In other words, $\partial_{\beta_{1}}f<0$ if $\beta_{0}>\beta_{0}^{*}$ and $\partial_{\beta_{1}}f>0$ if $\beta_{0}<\beta_{0}^{*}$. ∎
Lemma 10.
@@
where the last inequality is guaranteed by the constraints on $c$, as shown in Lemma 8. Meanwhile,
@@
where the right-hand side is a positive bound by the constraints on $c$ as shown in Lemma 8. ∎
Now we are ready to present the complete result on the 1-dimensional interval.
Theorem 11.
@@
Now we consider the case $\beta_{0}\beta_{1}>\lambda$. From (19), it is obvious that $g$ is monotone decreasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46). We now need to check whether $a^{*}\in[0,1-c]$. Notice that $a^{*}>1-c$ if and only if
@@
Similarly we consider the case $\beta_{0}\beta_{1}<\lambda$. From (19), it is obvious that $g$ is monotone increasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46). We now need to check whether $a^{*}\in[0,1-c]$. Notice that $a^{*}>1-c$ if and only if
@@
This completes the proof together with Proposition 9. ∎
Remark on homogeneous Robin boundary

### v2

# One-dimensional optimisation of indefinite-weight principal eigenvalues under inhomogeneous Robin conditions and a Schrödinger-type perturbation

@@ Abstract
We study the minimisation of the positive principal eigenvalue for an indefinite-weight problem under inhomogeneous Robin boundary conditions. The model is motivated by diffusive logistic equations in spatially heterogeneous environments, where the weight describes allocatable favourable resources and the Robin parameters measure boundary loss. After recalling the variational setting and the bang–bang reduction, we solve the one-dimensional optimisation problem completely: the optimal favourable set is an interval, and we determine whether it attaches to the left endpoint, to the right endpoint, or lies in the interior according to explicit criteria involving the two boundary parameters. The classification is then verified numerically by a transfer–matrix shooting method. We also introduce a Schrödinger-type extension with a fixed nonnegative background potential. In the coercive case we establish the corresponding principal-eigenvalue and bang–bang results, and in one dimension with constant potential we prove a stability result showing that minimisers for small background potential converge to those of the unperturbed problem. This perturbed regime is likewise illustrated numerically.
Keywords: principal eigenvalue; indefinite weight; Robin boundary condition; shape optimisation; transfer–matrix method; Schrödinger-type perturbation
MSC Classification: 35P15, 35J25, 49Q10, 92D25
@@ 1 Introduction
Spatial heterogeneity, diffusion, and boundary exchange are basic features of many reaction–diffusion models arising in population dynamics, spatial ecology, and related applied fields [24, 21, 20, 7, 13]. In such models, dispersal is described by diffusion, local growth or decay depends on the position $x\in\Omega$, and the interaction with the exterior is encoded through boundary conditions. A standard prototype is the diffusive logistic equation on a bounded habitat,
@@
$\begin{cases}\displaystyle u_{t}=\Delta u+\omega u[m(x)-u]&\text{in }\Omega\times\mathbb{R}_{+},\\[1.72218pt] \partial_{\mathbf{n}}u+\beta_{0}u=0&\text{on }\partial\Omega\times\mathbb{R}_{+},\\[1.72218pt] u(x,0)\geq 0,\quad u(x,0)\not\equiv 0&\text{in }\overline{\Omega},\end{cases}$ (1)
where $\omega>0$ measures the overall reproductive intensity, $m$ is the intrinsic growth rate, and $\beta_{0}$ measures the inhospitableness of the boundary. When $\beta_{0}=0$ the boundary is impermeable, while positive values of $\beta_{0}$ model leakage, exchange, or mortality at the boundary; such Robin conditions are also standard in broader diffusion, thermal, and quantum models [8, 11, 19, 9]. In a heterogeneous habitat, the sign-changing coefficient $m$ distinguishes favourable and unfavourable regions, and the principal eigenvalue of the linearised operator determines the threshold between extinction and persistence [6, 25, 3, 18].
This leads to the indefinite-weight eigenvalue problem
$\left\{\begin{aligned} \Delta\phi+\lambda m\phi&=0&&\text{in }\Omega,\\ \partial_{\mathbf{n}}\phi+\beta\phi&=0&&\text{on }\partial\Omega,\end{aligned}\right$. (2)
where $\Omega\subset\mathbb{R}^{n}$ is a bounded Lipschitz domain, $\mathbf{n}$ is the unit outer normal on $\partial\Omega$, the weight $m\in L^{\infty}(\Omega)$ changes sign in $\Omega$, and
$1\leq m(x)\leq\kappa\qquad\text{for a.e. }x\in\Omega,$
for some fixed constant $\kappa>0$. Throughout the paper, $\beta:\partial\Omega\to\mathbb{R}$ is a piecewise continuous Robin parameter. Following [5, 23, 1, 12], a real number $\lambda$ is called a principal eigenvalue of (2) if the associated eigenfunction $\phi\in H^{1}(\Omega)$ is positive.
The optimisation problem appears when the total amount of favourable habitat is limited. One imposes pointwise bounds on $m$ together with an integral constraint on the total resource,
$\int_{\Omega}m\leq-m_{0}|\Omega|,$
with $m_{0}$ fixed, and then asks how the favourable part of the habitat should be placed so as to minimise the positive principal eigenvalue. From the modelling point of view, the amount of favourable resource is prescribed, but its spatial arrangement is free; lowering the principal eigenvalue lowers the persistence threshold and therefore enlarges the range of viable reproduction parameters. Questions of this type belong to the broader spectral-optimisation tradition going back at least to Bôcher and Kreĭn [4, 16]. In the indefinite-weight setting, the modern variational theory was developed for Neumann boundary conditions by Lou and Yanagida [18], for Robin boundary conditions by Hintermüller, Kao and Laurain [14], and further analysed by Lamboley, Laurain, Nadin and Privat [17]. Beyond the interval case, complete structural results remain rare; the best understood higher-dimensional examples include cylindrical domains [15] and spherical shells [22].
The first aim of the present paper is to treat the problem under inhomogeneous Robin boundary conditions in one dimension, allowing distinct nonnegative parameters at the two endpoints. This asymmetric setting is natural when the two ends of the habitat interact differently with the exterior. Our main theorem gives a complete classification of the optimal favourable interval: depending on the boundary parameters and the corresponding eigenvalue, the optimal interval may attach to the left endpoint, attach to the right endpoint, or lie in the interior. We then verify this classification numerically by a transfer–matrix shooting method.
The second aim is to record a natural Schrödinger-type extension in which the optimisable indefinite weight is accompanied by a fixed hostile background potential. In ecological terms, this additional potential may represent mortality, harvesting, pollution, predation, or some other adverse effect that is not itself subject to optimisation [10]. In the coercive case we show that the principal-eigenvalue theory and the bang–bang structure remain valid. We then study the one-dimensional family with constant potential and prove a stability result showing that minimisers for small positive potential converge to those of the unperturbed problem. A second numerical study, again based on transfer matrices, confirms this perturbative picture for representative canonical regimes.
The paper is organised as follows. After the variational preliminaries and the bang–bang reduction in the remainder of this section, Section 2 contains the complete one-dimensional analysis for the original indefinite-weight problem. Section 3 introduces the Schrödinger-type extension, first at the general variational level and then for the one-dimensional constant-potential family. Section 4 presents the two numerical verification procedures: first for the unperturbed one-dimensional classification, and then for the perturbative constant-potential regime. We conclude with a brief discussion of the analytical scope of the results, their modelling interpretation, and possible directions for further work.
@@
We first prove the existence and uniqueness of a positive principal eigenvalue following the method in [1].
Lemma 1.1 ([25], Theorem 11.4 and 11.10).
@@
the map $\mu:\lambda\mapsto\mu(\lambda)$ is also concave as the infimum of a collection of concave functions. In other words, $\mu^{\prime\prime}(\lambda)\leq 0$. Meanwhile, it is clear that there exist $\phi_{1},\phi_{2}\in H^{1}(\Omega)$ satisfying $\int_{\Omega}m\phi_{1}^{2}>0$ and $\int_{\Omega}m\phi_{2}^{2}<0$, hence $\lim_{\lambda\to+\infty}\mu(\lambda)=\lim_{\lambda\to-\infty}\mu(\lambda)=$-$\infty$. Thus, $\mu=\mu(\lambda)$ increases until it attains its maximum, and decreases thereafter.
@@
and thus $\mu(0)=0$. Since the associated eigenfunction is clearly a constant, by Lemma 1.2, we get:
@@
  • if $\int_{\Omega}m=0$, then $\mu^{\prime}(0)=0$, and $\mu$ has no other zeros, hence (2) has a unique principal eigenvalue 0.
@@ 2 The 1-dimensional case
In this section, we solve problem (2) completely in one dimension.
@@
Now let the optimal set be the interval $E=(a,b)$ where $0<a<b<1$ and c=bac=b-a. We determine the optimal value of $a$ by adapting the method of [14]. Let $\beta_{0}=\beta(0)\geq 0$ and $\beta_{1}=\beta(1)>0$. Rewrite the problem using $m=\kappa\chi_{E}-\chi_{\Omega\setminus E}$, and we have
@@
where $C_{i}=C_{i}(E)$ are constants for $1\leq i\leq 6$. We now simplify the boundary and transmission conditions in (6). Collecting the coefficients of $C_{i}$ ($1\leq i\leq 6$), we obtain
@@
Rearranging the terms gives
@@
It is useful to check (11) in two extreme cases. When $\beta_{0}=\beta_{1}=0$, the problem reduces to the Neumann case, and we obtain the characteristic equation [^5] We remark that the characteristic equations (18) and (20) obtained in [14] based on homogeneous Robin boundary condition admit the same typo on the left-hand side.
@@
On the other hand, when $\beta_{0},\beta_{1}\to+\infty$, the problem reduces to the Dirichlet case, and we similarly obtain the characteristic equation
@@
We next use the product-to-sum formulae for hyperbolic functions to simplify this expression. Using $a\sqrt{\lambda}+\sqrt{\lambda}(1-b)=\sqrt{\lambda}(1-c)$ we get
@@
It is shown in Theorem 1.4 that $\partial_{\beta_{0}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$ and $\partial_{\beta_{1}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$. Now we have the implications:
@@
We prove in Lemma 2.2 that $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)$ cannot vanish for any $a\in[0,1-c]$. Therefore, for each $a$, it has a unique zero
@@
Denote by $h=h(c,\kappa,\lambda)$ the expression on the right-hand side of (39). We show in Lemma 2.3 that $h<1$ for any $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$ under the assumed constraints on $c$, so $A(a)<0$ for any $a\in[0,1-c]$. In other words, $\partial_{\beta_{1}}f<0$ if $\beta_{0}>\beta_{0}^{*}$ and $\partial_{\beta_{1}}f>0$ if $\beta_{0}<\beta_{0}^{*}$. ∎
Lemma 2.5.
@@
where the last inequality is guaranteed by the constraints on $c$, as shown in Lemma 2.3. Meanwhile,
@@
where the right-hand side is a positive bound by the constraints on $c$ as shown in Lemma 2.3. ∎
We are now in a position to state the complete result for the one-dimensional interval.
Theorem 2.6.
@@
Now we consider the case $\beta_{0}\beta_{1}>\lambda$. From (19), it is clear that $g$ is monotone decreasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46). We now need to check whether $a^{*}\in[0,1-c]$. Notice that $a^{*}>1-c$ if and only if
@@
Similarly we consider the case $\beta_{0}\beta_{1}<\lambda$. From (19), it is clear that $g$ is monotone increasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46). We now need to check whether $a^{*}\in[0,1-c]$. Notice that $a^{*}>1-c$ if and only if
@@
This completes the proof together with Proposition 2.4. ∎
Remark on homogeneous Robin boundary.